OpenStudy (anonymous):

What is the surface area of a sphere with radius of 3.5? A. 12.25p B. 14p C. 24.5p D. 49p

OpenStudy (anonymous):

do you know formula for calculating surface area of a sphere?

OpenStudy (anonymous):

yes but I keep getting 153.94

OpenStudy (anonymous):

In all answers there are p, so I guess its \(\pi\)? So you should not write 3.14 instead of it but just keep it.

OpenStudy (anonymous):

this is due today and I have no idea what you just said

OpenStudy (anonymous):

write down formula of sphere for me

OpenStudy (anonymous):

4 pi r 2

OpenStudy (anonymous):

ok so \[S=4\pi r^2=4\pi3.5^2=4\pi12.25=49\pi\] and taht's your answer, its D you don't need to use 3.14 instead of \(\pi\) because no one asked you. 3.14 is just approximation of \(\pi\) and for math you need the most accurate result.

OpenStudy (anonymous):

thanks but I found it on yahoo answers

OpenStudy (anonymous):

do you understand why its not 153.94?

OpenStudy (anonymous):

no but who cares